OpenStudy (anonymous):

2ne^(-n) can we use the integral test to test for its convergence?

OpenStudy (anonymous):

Its supposed to be from 0 to infinity! ;/

OpenStudy (tkhunny):

Parts? \(\int 2xe^{-x}\;dx\;=\;\int 2x\;d\left(-e^{-x}\right) = -2xe^{-x} + \int e^{-x}\;d(2x)\;=\;-2xe^{-x} + 2\int e^{-x}\;dx\)